Istituto ed Osservatorio Astronomico Referred and Invited Papers 1997-1993

(written Jul 2 1997)

[1] I. Pagano, R. Ventura, M. Rodono, G. Peres & G. Micela, “A major optical flare on the recently discovered X-ray active dMe star G102-21,” A&A 318 (1997), 467.

[2] F. Favata, A. Maggio, G. Peres & S. Sciortino, “On the Coronal Abundances Derived with the SAX LEX and ASCA SIS detectors ,” A&A in press (1997).

[3] A. Ortolani, A. Maggio, R. Pallavicini, S. Sciortino, J.J. Drake & S.A. Drake, “The X-ray Coronal Emission of Lambda Andromedae Observed with ASCA and ROSAT,” A&A in press (1997).

[4] S. Orlando, Y.-Q. Lou, G. Peres & R. Rosner, “Alfvenic fluctuations in fast and slow solar winds ,” J. Geophys. Res. in press (1997).

[5] F. Bocchino, A. Maggio & S. Sciortino, “ROSAT PSPC Observation of the North-East Region of the Vela SNR: II. Spectral Analysis with a Non-Equilibrium of Ionization Emission Model,” ApJ 481 (1997), 872.

[6] A. Maggio & G. Peres, “Loop Modeling of Coronal X-ray Spectra. IV. One- and Two-loop model fitting of ROSAT/PSPC spectra: three test cases,” A&A in press (1997).

[7] F. Reale, G. Peres & S. Serio, “Radiatively driven downdrafts and redshifts in transition region lines: - II.Exploring the parameter space ,” A&A 318 (1997), 506.

[8] R. Betta, G. Peres, F. Reale & S. Serio, “An Adaptive Grid Code for High Resolution 1-D Hydrodynamics of the Solar and Stellar Transition Region and Corona,” A&AS 122 (1997), 585.

[9] A. Ciaravella, A. Maggio & G. Peres, “Loop Modeling of Coronal X-ray Spectra: III - Fitting loop spectra with one- and two-components thermal models,” A&A in press (1997).

[10] G. Micela, J. Pye & S. Sciortino, “Coronal Properties of Old Disk and Halo Nearby dM Stars,” A&A 320 (1997), 865 .

[11] F. Damiani, A. Maggio, G. Micela & S. Sciortino, “A Method Based on Wavelet Transforms for Source Detection in Photon-Counting-Detector Images. I: Theory and General Properties ,” ApJ 483 (1997), 350.

[12] F. Damiani, A. Maggio, G. Micela & S. Sciortino, “A Method Based on Wavelet Transforms for Source Detection in Photon-Counting- Detector Images. II: Application to ROSAT PSPC Images ,” ApJ 483 (1997), 370.

[13] M. Barbera, A. Collura, A. Dara, M. Leone, F.R. Powell, S. Serio, S. Varisco & M.V. Zombeck, “Effects of interference and oxidation on the UV/visible rejection properties of filters for soft X-ray detectors,” Experimental Astronomy 7 (1997), 51–63.

[14] F. Reale, R. Betta, G. Peres, S. Serio & J. McTiernan, “Determination of the length of coronal loops from the decay of X-ray flares: I. Solar flares observed with Yohkoh SXT,” A&A in press (1997).

[15] F. Favata, G. Micela & S. Sciortino, “On the ubiquitous Weak-Line T-Tauri population detected in the ROSAT All-Sky Survey,” A&A in press (1997).

[16] G. Micela, F. Favata & S. Sciortino, “Hipparcos distances of X-ray selected stars: implications on their nature as stellar population,” A&A in press (1997).

[17] F. Favata, R. Mewe, N. S. Brickhouse, R. Pallavicini & G. Micela, “A SAX/LECS x-ray observation of the active binary Capella,” A&A in press (1997).

[18] M.V. Zombeck, M. Barbera, A. Collura & S.S. Murray, “An explanation of the ROSAT HRI UV sensitivity,” ApJ Letters in press (1997).

[19] A. Ortolani, R. Pallavicini & G. Tagliaferri, “X-ray Spectroscopy of RS CVn Binaries: the Quiescent and Flaring EXOSAT Spectra Revisited,” A&A in press (1997).

[20] G. Peres, S. Orlando, F. Reale, R. Rosner & H. Hudson, “The Sun as an X-Ray Star: Overview of the Method. Implication of Yohkoh data for the solar-stellar connection,” in Observational Plasma Astrophysics: Five Years of YOHKOH and beyond: Proceedings of the Yohkoh 5th Anniversary Symposium, T. Watanabe, T. Kosugi & A. C. Sterling, eds., Kluwer Academic Publishers, 1997, in press, Invited talk presented at the Yohkoh 5th Anniversary Symposium, held in November 1996, in Yoyogi, Tokyo, Japan.

[21] A. Maggio, G. Micela & G. Peres, “Loop modeling of x-ray coronae: solartype stars in young clusters and associations,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997 .

[22] J. F. Schachter, R. Remillard, S. Saar, F. Favata, S. Sciortino & M. Barbera, “Optical and X-ray Characteristics of Stars Detected in the Einstein Slew Survey,” Ap.J. 463 (1996), 747–765.

[23] G. Micela, S. Sciortino, V. Kashyap, F. R. Harnden, Jr. & R. Rosner, “ROSAT Observations of the Pleiades Cluster. I - X-ray Characteristcs of a Coeval Stellar Population,” ApJS 102 (1996), 75.

[24] A. Ciaravella, G. Peres, A. Maggio & S. Serio, “Loop Modeling of Coronal X-ray Spectra: I - general properties,” A&A 306 (1996), 553.

[25] A. Maggio & G. Peres, “Loop Modeling of Coronal X-ray Spectra: II - Parameter Constraints and Diagnostics,” A&A 306 (1996), 563.

[26] F. Reale, G. Peres & S. Serio, “Radiatively driven downdrafts and redshifts in transition region lines: - I. Reference model,” A&A 316 (1996), 215.

[27] F. Favata, G. Micela & S. Sciortino, “Lithium abundance in a volumelimited sample of nearby main sequence G and K stars,” A&A 311 (1996), 951–960.

[28] G. Cutispoto, G. Tagliaferri, M. Rodono‘, R. Pallavicini & L. Pasquini, “Photometric and Spectroscopic Studies of Cool Stars Discovered in EX-OSAT X-ray Images. III. Photometric Properties,” Astronomy and Astrophysics 115 (1996), 41.

[29] S. Randich, N. Aharpour, R. Pallavicini, C.F. Prosser & J.R. Stauffer, “Lithium Abundance in the Young Open Cluster IC 2602 ,” A&A in press (1996).

[30] F. Morale, G. Micela, F. Favata & S. Sciortino, “Stroemgren four-colours Photometry of X-Ray Active Late-Type Stars: Evidence for Activityinduced Deficiency in the m1 Index.,” A&AS 119 (1996), 403 .

[31] L. Pasquini, S. Randich & R. Pallavicini, “Lithium in M67: Evidence for Spread in a Solar Age Cluster,” A&A Submitted (1996).

[32] R. Mewe, J. S. Kaastra, S. M. White & R. Pallavicini, “Simultaneous EUVE and ASCA Observations of AB Dor: Temperature Structure and Abundances of the Quiescent Corona,” A&A 315 (1996), 170.

[33] F. Favata, G. Micela & S. Sciortino, “The Relationship Between Lithium and Activity in Disk Population Main Sequence G and K stars,” A&A in press (1996).

[34] G.E. Villa, A.F. Abbey, M. Arnaud, M. Balasini, G.F. Bignami, M. Boer, H. Brauninger, I. Butler, G. Cafagna, C. Cara, C. Cahabaud, L. Chiappetti, R. Cole, M. Conte, J. Dowson, P. Dhez, R. Duc, G. Di Cocco, P. Ferrando, C.V. Goodall, H. Hippman, A. Holland, E. Kendziorra, A. Labeque, N. Meidinger, C. Moriggio, N. La Palombara, C. Musso, E. Poindron, G. Peres, E. Pferfferman, C. Pigot, J.P. Pye, C. Reppin, D. Schmitt, S. Sciortino, S. Serio, J. Spragg, J.B. Stephen, L. Struder, M. Trifoglio, P. Tua, M.J.L. Turner & S. Whitehead, “The EPIC system onboard the ESA XMM mission,” in SPIE Proceedings #2808, 1996, 402, Presented at the SPIE’s Annual Meeting, Denver (Colorado), U.S.A., 4-9 August 1996.

[35] A. Collura, M. Barbera, A. Dara, G. Inzerillo, F. Mirabello, S. Sciortino & S. Serio, “Calibration of the AXAF-HRC UV/Ion shields at Osservatorio Astronomico di Palermo G.S.Vaiana: I - Instrumental set-up ,” in SPIE Proceedings #2808, 1996, 144, Presented at the SPIE’s Annual Meeting, Denver (Colorado), U.S.A., 4-9 August 1996 .

[36] A. Collura, M. Barbera, A. Dara & M.V. Zombeck, “Calibration of the AXAF-HRC UV/Ion shields at Osservatorio Astronomico di Palermo G.S.Vaiana: II - X-ray transmission measurements,” in SPIE Proceedings #2808, 1996, 134, Presented at the SPIE’s Annual Meeting, Denver (Colorado), U.S.A., 4-9 August 1996 .

[37] M. Barbera, A. Collura & A. Dara, “Calibration of the AXAF-HRC UV/Ion shields at Osservatorio Astronomico di Palermo G.S.Vaiana: III

- Synchrotron measurements of XANES in aluminum coated Lexan film samples ,” in SPIE Proceedings #2808, 1996, 120, Presented at the SPIE’s Annual Meeting, Denver (Colorado), U.S.A., 4-9 August 1996 .

[38] M. Barbera, A. Collura, A. Dara & M.V. Zombeck, “Calibration of the AXAF-HRC UV/Ion shields at Osservatorio Astronomico di Palermo G.S.Vaiana: IV - UV rejection measurements ,” in SPIE Proceedings #2808, 1996, 108, Presented at the SPIE’s Annual Meeting, Denver (Colorado), U.S.A., 4-9 August 1996 .

[39] G.R. Meehan, A.T. Kenter, R.P. Kraft, S.S. Murray, M.V. Zombeck, K. Kobayashi, J. Chappell, M. Barbera & A. Collura, “Measurement of the transmission of the UV/ion shields for the AXAF high-resolution camera,” in SPIE Proceedings #2808, 1996, 210, Presented at the SPIE’s Annual Meeting, Denver (Colorado), U.S.A., 4-9 August 1996 .

[40] F. Favata, G. Micela & S. Sciortino, “The [Fe/H] distribution of a volume limited sample of solar-type stars and its implications for Galactic chemical evolution,” A&A in press (1996).

[41] S. Orlando, Y.-Q. Lou, R. Rosner & G. Peres, “Propagation of threedimensional Alfven waves in stratified, thermally conducting solar wind,” J. Geophys. Res. 101 (1996), 24443.

[42] F. Favata, G. Micela, S. Sciortino & F. Morale, “On the comparison between spectroscopic and photometric metallicity measurements in active solar-type stars,” A&A in press (1996).

[43] G. Tagliaferri, S. Covino, T.A. Fleming, M. Gagne’, R. Pallavicini, F. Haardt & Y. Uchida, “Coronal X-ray emission of HD 35850: the ASCA view,” A&A in press (1996).

[44] S. Serio, “Evidence of Magnetic Coronae in other Stars ,” in Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity, Y. Uchida, T. Kosugi & H.S. Hudson, eds., Kluwer Academic Publishers, Dordrecht, The Netherlands, 1996, 63 , Presented at the IAU Colloquium 153 - Magnetodynamic Phenomena in the Solar Atmosphere

- Prototypes of Stellar Activity - Tokyo (Japan) 22-26 May 1995.

[45] G. Peres & S. Orlando, “Structure and Stability of EUV Loops Originating from Sunspots,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Series, 1996 , 93, Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[46] G. Micela, “X-ray Observations of Open Clusters,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Series , 1996 , 347 , Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[47] R. Pallavicini, “Spectroscopic Diagnostics of Astrophysical Plasmas,” in Plasma Astrophysics, C. Chiuderi & L. Einaudi, eds., Lecture Notes in Physics, 1996, 241.

[48] R. Pallavicini, “Stellar Magnetic Activity, Activity Cycles, and Dynamos,” in Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Magnetic Activity, Y. Uchida, T. Kosugi & H.S. Hudson, eds., Kluwer Academic Publishers, Dordrecht, The Netherlands, 1996 , 359, Presented at the IAU Colloquium 153 - Magnetodynamic Phenomena in the Solar A tmosphere - Prototypes of Stellar Activity - Tokyo (Japan) 22-26 May 1995.

[49] R. Pallavicini & S.M. White, “Millimeter Continuum Observations of Stars,” in Science with large Millimetre Arrays, P.A. Shaver, ed., 1996 , 268, Presented at ESO-IRAM-NFRA-Onsala Workshop on Science with Large Millimetre Arrays, Garching bei Munchen, 11-13 December 1995 .

[50] I. Chinnici, “La voyage scientifique de G. Rayet en Italie et la fondation de L’Observatoire de Bordeaux,” 1996, Presented at the 121st Congres National des Societes historiques et scientifiques, Nice, 26-31 octobre 1996.

[51] I. Chinnici, “An ”italian” Observatory in India: the history of the Calcutta Observatory,” Studies in History of Medicine and Science 14 (1996).

[52] R. Pallavicini, “Prospects for Stellar Coronae Observations,” in The Next Generation of X-Ray Observatories, M. Turner, ed., 1996, in press.

[53] S. Sciortino, “Stellar Coronal Observations,” 1996, Invited Presentation at the XXX Congress of the S.A.It - May 1997, Monte Porzio.

[54] G. Foderà Serio, “Giuseppe Tomasi di Lampedusa nell’astronomia,” 1996, Relazione su invito al Convegno ”Il Gattopardo oltre la letteratura”, Roma, 5-6 Dicembre 1996.

[55] I. Chinnici, “Gli Strumenti Astronomici del Principe Tomasi di Lampedusa,” 1996, Relazione su invito al Convegno ”Il Gattopardo oltre la letteratura”, Roma, 5-6 Dicembre 1996.

[56] G. Foderà Serio, “Il Gattopardo e l’astronomia,” 1996, Presentazione su invito della serie ”Il Meridione e le Scienze: Personaggi ed Interpreti”, Napoli.

[57] F. Damiani, G. Micela, S. Sciortino & F. R. Harnden, Jr., “Einstein Observations of T-Tauri Stars in Taurus-Auriga. I- Properties of X-ray Emission,” Ap. J. 446 (1995), 331.

[58] F. Damiani & G. Micela, “Einstein Observations of T-Tauri Stars in Taurus-Auriga. II - Relationships between X-ray Emission and Pre-Main-Sequence Activity,” Ap. J. 446 (1995), 341.

[59] S. Sciortino, F. Favata & G. Micela, “The Stellar Coronal Component of the Galaxy: II - An Analysis of the Stellar Content of the Einstein Extended Medium Sensitivity Survey.,” Astron. Astrophy. 296 (1995), 370.

[60] A. Collura, G. Micela, S. Sciortino, F. R. Harnden, Jr. & R. Rosner, “An Objective Multicolor Method for the Characterization of Low Resolution X-ray Spectra,” ApJ 446 (1995), 108.

[61] S. Orlando, G. Peres & S. Serio, “Models of stationary siphon flows in stratified, thermally conducting coronal loops: I - regular solutions,” Astr. Ap. 294 (1995), 861.

[62] F. Reale & G. Peres, “Solar Flare X-ray Imaging: coronal loop hydrodynamics and diagnostics of the rising phase,” Astr. Ap. 299 (1995), 225.

[63] F. Reale, “Thermal Conduction in a 2-D FCT Plasma Hydrodynamic Code ,” Computer Phys. Comm. 86 (1995), 13.

[64] F. Favata, M. Barbera, G. Micela & S. Sciortino, “Lithium, X-ray Activity and Rotation in an X-ray Selected Sample of Solar-Type Stars,” A&A 295 (1995), 147 .

[65] G. Micela, F. Favata, J. Pye & S. Sciortino, “The Discovery of a Very Active High Proper Motion dMe Binary Star,” A&A 298 (1995), 505.

[66] F. Favata, G. Micela & S. Sciortino, “The Space Density of Active Binaries from X-ray Surveys,” A&A 298 (1995), 482.

[67] F. Favata, G. Micela & S. Sciortino, “High Resolution Spectroscopy of Old Late K Dwarf Stars Around the Lithium I 6707.8 A line: Is Lithium There ?,” A&AL 297 (1995), L1.

[68] A. Maggio, S. Sciortino, A. Collura & F. R. Harnden, Jr., “ROSAT PSPC Spectral Fitting Simulations with One- and Two-Temperature Models of Optical-Thin Plasmas,” A&AS 110 (1995), 573.

[69] S. Orlando, G. Peres & S. Serio, “Models of stationary siphon flows in stratified, thermally conducting coronal loops: II - Shocked solutions,” Astr. Ap. 300 (1995), 549.

[70] D. Spadaro, S. Orlando, G. Peres & P. Leto, “Effects of flows and nonequilibrium ionization on some C IV line ratios on the Sun,” Astr. Ap. 302 (1995), 285.

[71] R. Ventura, G. Peres, I. Pagano & M. Rodono, “Observations of small flares and a long enduring pre-flare dip on V1054 Oph,” A&A 303 (1995), 509.

[72] G. Cutispoto, R. Pallavicini, M. Rodono‘ & M. Kurster, “Photometry of Cool Stars detected in Extreme-Ultraviolet (EUV) All-Sky Surveys,” Astronomy and Astrophysics 297 (1995), 770.

[73] F. Reale, “Metodi numerici in idrodinamica,” Atti Accademia Peloritana dei Pericolanti, Classe I di Scienze Fis. Mat. e Nat. LXXII, Suppl. N.1 (1995), in press, Scuola Nazionale di Astrofisica, ciclo 93-94: IV corso, Astrofisica dei plasmi - galassie, Palermo, 17-21 ottobre 1994.

[74] I. Chinnici, “Il Contributo di Padre F. Denza alla Ricerca Astronomica ,” 1995, Invited Paper al convegno “Padre Francesco Denza nel Centenario della Morte”, Moncalieri 14-15 Dicembre 1994.

[75] R. Pallavicini, “Flares in Late-Type Stars: X-ray,” in Flares and Flashes, J. Greiner, H. Duerbeck & R. Gershberg, eds., 1995, 148.

[76] R. Pallavicini, “Stellar Magnetic Activity, Activity Cycles, and Dynamos,” in Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Magnetic Activity, Y. Uchida, T. Kosugi & H.S. Hudson, eds., Kluwer Academic Publishers, Dordrecht, The Netherlands, 1996 , 359, Presented at the IAU Colloquium 153 - Magnetodynamic Phenomena in the Solar A tmosphere - Prototypes of Stellar Activity - Tokyo (Japan) 22-26 May 1995.

[77] V. Kashyap, R. Rosner, F. R. Harnden, Jr., A. Maggio, G. Micela & S. Sciortino, “X-ray Emission on Hybrid Stars: Rosat Observation of alpha Tra and iota Aur,” Ap. J. 431 (1994), 402.

[78] A. Maggio, S. Sciortino & F. R. Harnden, Jr., “ROSAT X-ray Observations of Late-Type Evolved Stars: On The Relationship between Coronal Temperatures and Luminosities.,” Ap. J. 432 (1994), 701.

[79] G. Peres, F. Reale & L. Golub, “Loop Models of Low Coronal Structures Observed by NIXT,” Ap. J. 422 (1994), 412.

[80] A. Collura, F. Reale, E. Schulman & J. N. Bregman, “Two Very Luminous X-ray Sources in M82,” Ap. J. (Letters) 420 (1994), L63.

[81] A. Maggio, F. Reale, G. Peres & A. Ciaravella, “The Analysis System for Astrophysical Plasmas (ASAP) of the Osservatorio Astronomico di Palermo,” Comp. Phys. Comm. 81 (1994), 105.

[82] F. Bocchino, A. Maggio & S. Sciortino, “ROSAT PSPC Observation of the North-East Region of the Vela SNR: I. Evidence of Thermal Structure on Arcminutes Scale.,” ApJ 437 (1994), 209.

[83] B. Schmieder, G. Peres, S. Enome, R. Falciani, P. Heinzel, J. C. Hénoux, J. Mariska, F. Reale, M. L. Rilee, B. Rompolt, K. Shibasaki, A. V. Stepanov, J. P. Wuelser, D. Zarro & V. Zharkova, “Energy Transport and Dynamics,” Solar Physics 153 (1994), 55, Results of Team 4 of the Flare 22 workshop meeting held in Ottawa.

[84] F. Reale, S. Serio & G. Peres, “Thermally Unstable Perturbations in Stratified Conducting Atmospheres,” ApJ 433 (1994), 811.

[85] F. Reale, F. Bocchino & S. Sciortino, “Parallel Computing on Unix Workstation Arrays,” Computer Phys. Comm. 83 (1994), 130.

[86] F. Damiani, G. Micela, S. Sciortino & F. R. Harnden, Jr., “Einstein X-ray Observations of Herbig Ae/Be Stars,” ApJ 436 (1994), 807.

[87] M. Barbera, G. Austin, A. Collura, K. Flanagan, S. Jelinsky, S. Murray, S. Serio & M. Zombeck, “Development of the UV/Ion shields for the AXAF HRC,” in SPIE Prooceedings Vol. 2280, Siegmund O. H. W. & Vallerga J. V., eds., 1994, 214, Presented at the SPIE’s Annual Meeting, San Diego, California, July 27-28.

[88] M. Barbera, A. Collura, F. Favata, G. Inzerillo & F. Mirabello, “Simple Gas Scintillation Proportional Counter soft X-ray detector for laboratory usage,” in SPIE Prooceedings Vol. 2280, Siegmund O. H. W. & Vallerga J. V., eds., 1994, 154, Presented at the SPIE’s Annual Meeting, San Diego, California, July 27-28.

[89] A. Collura, M. Barbera, G. Inzerillo, F. Mirabello, S. Sciortino & S. Serio, “X-ray Astronomy Calibration and Testing (XACT) Facility at the G.S. Vaiana Observatory in Palermo,” in SPIE Prooceedings Vol. 2280, Siegmund O. H. W. & Vallerga J. V., eds., 1994, 206, Presented at the SPIE’s Annual Meeting, San Diego, California, July 27-28.

[90] M. Barbera, D. Breslau, K. Flanagan, D. Graessle & M. Zombeck, “Synchrotron X-ray Transmission Measurements in the Calibration Program for the UV/Ion Shields of the AXAF HRC,” in SPIE Prooceedings Vol. 2280, Siegmund O. H. W. & Vallerga J. V., eds., 1994, 229, Presented at the SPIE’s Annual Meeting, San Diego, California, July 27-28.

[91] K. Flanagan, M. Barbera, S. Murray & M. Zombeck, “Calibration Program for the AXAF High Resolution Camera,” in SPIE Prooceedings Vol. 2280, Siegmund O. H. W. & Vallerga J. V., eds., 1994, 243, Presented at the SPIE’s Annual Meeting, San Diego, California, July 27-28.

[92] R.J. Aucoin, T.H. Markert, S.A. Nenonen, M. Barbera & K.A. Flanagan, “Thin Polyimide and Silicon Nitrade Windows for Gas Proportional Counters,” in SPIE Prooceedings Vol. 2280, Siegmund O. H. W. & Vallerga J. V., eds., 1994, 134, Presented at the SPIE’s Annual Meeting, San Diego, California, July 27-28.

[93] G. Foderà, F. Martines & D. Randazzo, “Cataloghi di Strumenti Scientifici nella Biblioteca dell’Osservatorio Astronomico di Palermo,” Nuncius 2 (1994), 759.

[94] I. Chinnici, “Il Fondo “Tacchini” dell’Ufficio Centrale di Ecologia Agraria in Roma,” Quaderni P.RI.ST.EM. 7 (1994).

[95] G.F. Gahm, H. Zinnecker, R. Pallavicini & L. Pasquini, “A Search for Cold Dust around Post-T Tauri Stars,” A&A 282 (1994), 123.

[96] S. Randich, M.S. Giampapa & R. Pallavicini, “Lithium in RS CVn Binaries and Related Chromospherically Active Stars. III. Northern RS CVn Systems,” A&A 283 (1994), 893.

[97] G. Tagliaferri, G. Cutispoto, R. Pallavicini, S. Randich & L. Pasquini, “Photometric and Spectroscopic Studies of Cool Stars Discovered in EX-OSAT X-ray Images. II. Lithium Abundances,” A&A 285 (1994), 272.

[98] L. Pasquini, Q. Liu & R. Pallavicini, “Lithium Abundances in Nearby Solar-like Stars,” Astronomy and Astrophysics 287 (1994), 205.

[99] S. Serio, “Modelling of Quiet Sun Coronal Structures,” in Solar Coronal Structures, V. Rusin, P. Heinzel & J.-C. Vial, eds., VEDA, 1994, 169, Proceedings of the IAU 144 Colloquium ”Solar Coronal Structures”, held in Tatranska Lomnica, Slovacchia, September 20-24, 1993.

[100] F. Reale, F. Bocchino & S. Sciortino, “Parallel Computations in Astrophysics on Unix Workstation Arrays,” IEEE (1994), Presented at the Massively Parallel Computing Systems 1994, May 1-6, Ischia.

[101] F. Reale, “Metodi numerici in idrodinamica,” Atti Accademia Peloritana dei Pericolanti, Classe I di Scienze Fis. Mat. e Nat. LXXII, Suppl. N.1 (1995), in press, Scuola Nazionale di Astrofisica, ciclo 93-94: IV corso, Astrofisica dei plasmi - galassie, Palermo, 17-21 ottobre 1994.

[102] R. Pallavicini, “Magnetic Activity in Stars,” Reports in Astronomy XVIIA (1994), 409.

[103] E. Antonucci, M. A. Dodero, R. Martin, G. Peres, F. Reale & S. Serio, “Simulations of the CaXIX Spectral Emission from a Flaring Solar Coronal Loop: II. Impulsive Heating by Accelerated Electrons,” Ap. J. 413 (1993), 786.

[104] M. Barbera, G. Micela, S. Sciortino, F. R. Harnden, Jr. & R. Rosner, “X-ray Emission at Low-Mass End: Results from an Extensive Einstein Observatory Survey,” Ap. J. 414 (1993), 846–866.

[105] A. Ciaravella, G. Peres & S. Serio, “Thermal Conduction and Modeling of Static Stellar Coronal Loops,” Sol. Phys. 154 (1993), 45.

[106] A. Collura, A. Maggio, G. Micela, S. Sciortino, F. R. Harnden, Jr. & R. Rosner, “ROSAT X-ray Detection of Epsilon Tau: Revisiting the Coronal and Transition Region Emission of the Hyades Giants,” Ap. J. 416 (1993), 204–207.

[107] F. Favata, M. Barbera, G. Micela & S. Sciortino, “A Search for Yellow Young Disk Population Stars among EMSS X-ray Sources by means of Lithium Abundance Determination.,” Astr. Ap. 277 (1993), 428.

[108] G. Micela, S. Sciortino & F. Favata, “Stellar Birthrate in the Galaxy: Constraints from X-ray flux-limited surveys.,” Ap. J. 412 (1993), 618.

[109] G. Peres & F. Reale, “The Importance of Plasma Viscosity on X-ray Line Diagnostics of Solar Flares,” Astr. Ap. 267 (1993), 566.

[110] G. Peres, R. Ventura, I. Pagano & M. Rodono‘, “Low Amplitude Variability and Transient Periodicity Before a Large Flare in FF And,” Astr. Ap. 278 (1993), 179.

[111] G. Peres & F. Reale, “Detectability of Chromospheric Evaporation Fronts in Solar Flares,” A&A Lett. 275 (1993), L13.

[112] F. Reale, M. Barbera & S. Sciortino, “A Parallel 2-D Hydrodynamic FOR-TRAN Code for Astrophysical Applications on Meiko Computing Surface,” Future Generation Computer Systems 9 (1993), 19, Presented at the 3rd Meiko User’s Conference, Manchester (UK), 9-10 April 1992.

[113] F. Reale, S. Serio & G. Peres, “Dynamics of the Decay of Confined Stellar X-Ray Flares,” Astr. Ap. 272 (1993), 486.

[114] B. Sylwester, J. Sylwester, S. Serio, F. Reale, R. D. Bentley & A. Fludra, “Dynamics of Flaring Loops: III. Interpretation of Flare Evolution in Emission Measure - Temperature Diagram,” Astr. Ap. 267 (1993), 586.

[115] E. Schulman, J. N. Bregman, A. Collura, F. Reale & G. Peres, “Accurate Period Determination of an Eclipsing Binary X-ray Source in M33,” Ap. J. (Letters) 418 (1993), L67.

[116] L. Stella, G. Chincarini, O. Citterio, P. Conconi, T. Maccacaro, G. Tagliaferri, G. Trincheri, A. Wolter, G. F. Bignami, F. Bocchino, A. Maggio, G. Micela, S. Sciortino, S. Serio, A. Collura, P. Giommi, L. Maraschi, R. Pallavicini, S. Pellegrini & G. Peres, “A 15-30 arcsec Resolution Replica X-ray Optics for AXAF-S,” SPIE 2011 (1993), 149.

[117] F. Ventura, G. Foderà Serio & M. Hoskins, “Possible Tally Stones at Mnajdra, Malta,” Journal for the History of Astronomy 24 (1993), 171.

[118] R. Pallavicini, G. Cutispoto, S. Randich & R. Gratton, “The Effects of Stellar Surface Activity on the Strength of the Lithium 6708 A Line ,” A&A 267 (1993), 145.

[119] S. Randich, R. Gratton & R. Pallavicini, “Lithium in RS CVn Binaries and Related Chromospherically Active Stars. II. Spectrum Synthesis Analysis,” A&A 273 (1993), 194.

[120] G. Peres, F. Reale & S. Serio, “Hydrodynamics and Diagnostics of Coronal Loops Subject to Dynamic Heating,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 151, Invited presentation at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26, 1992.

[121] F. Reale, “An Introduction to the World of Transputers,” Mem. S.A.It. in press (1993), Invited Lecture at the “School Astronet 1990”, Cagliari Ottobre 1-3.

[122] S. Sciortino, “Data Analysis Systems and Databases in Astronomy: The Case of the OAP-CFA Einstein Stellar X-ray Database.,” Mem. SAIt. 64 (1993), 125, Invited Talk XXXV Congresso SAIt, Torino, Maggio 1991.

[123] S. Sciortino, “Stellar Coronal Emission: What We Have Learned From Pre-ROSAT Observations,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 211, Invited Talk at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26.

[124] G. Foderà Serio, “On The History Of The Palermo Astronomical Observatory,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 21.

[125] R. Pallavicini, “Time Variability of Stellar X-ray Emission,” in Physics of Solar and Stellar Coronae, J.F. Linsky & S. Serio, eds., 1993, 237.